OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }*\frac{ 1 }{ 4 }=\frac{ 3 }{ 20}\] You can't simplify it, so the answer is simply 3/20 mile. Understand how I got it?

OpenStudy (anonymous):

yup